FInd the standard form of an ellipse with verticies (-5,-6) and (-5,8) and a minor axis of 6
Answer by lwsshak3(11628)

FInd the standard form of an ellipse with verticies (-5,-6) and (-5,8) and a minor axis of 6
**
This is an ellipse with vertical major axis.
Its standard form of equation: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k)=(x,y) coordinates of center
x-coordinate of center=-5
y-coordinate of center= (8-6)/2=1 (midpoint formula)
center: (-5,1)
length of vertical major axis=14 (-6 to 8)=2a
a=7
a^2=49
..
length of minor axis=6=2b
b=3
b^2=9
..
Equation of given ellipse:
(x+5)^2/9+(y-1)^2/49=1